# Almost All Vectorial Functions Have No Extended-Affine Symmetries

### v1

@@ Abstract
We prove that asymptotically almost all vectorial functions over finite fields have trivial extended-affine stabilizers. As a consequence, the number of EA-equivalence classes is asymptotically equal to the naive estimate, namely the total number of functions divided by the size of the EA-group, with vanishing relative error. Furthermore, we derive upper bounds on collision probabilities for both extended-affine and CCZ equivalences. For EA-equivalence, we leverage the trivial-stabilizer result to establish a matching lower bound, yielding a tight asymptotic formula that shows two independently sampled functions are EA-equivalent with super-exponentially small probability. The results validate random sampling strategies for cryptographic primitive design and show that functions with nontrivial EA-symmetries form an exponentially rare subset.
Keywords: Boolean functions, extended-affine equivalence, group action, orbit counting, asymptotic enumeration
@@ 1 Introduction
Both questions are intimately connected to the structure of EA-stabilizer groups. By the orbit-stabilizer theorem, the size of each equivalence class is inversely proportional to the stabilizer size, which directly impacts both the total number of classes and the collision probability. However, explicit computation of stabilizers is infeasible except for very small dimensions, as the function space grows doubly exponentially with the input dimension.
@@ 1.1 Our Contributions
Collectively, these results demonstrate that functions with nontrivial EA-symmetries form an exponentially rare subset. This complements existing approaches focused on specific cryptographic properties and provides a theoretical foundation for understanding the prevalence of symmetries in the space of vectorial functions.
@@ 1.2 Related Work
Previous work on symmetries of Boolean functions has primarily focused on simpler group actions. Clausen [3] proved that almost all Boolean functions (not vectorial) have no linear symmetries under the action of $\mathrm{GL}(n,2)$. Clote and Kranakis [4] established similar results for invariance groups under variable permutations. However, these results apply only to Boolean functions $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ and do not address EA-equivalence, which involves affine transformations on both domain and range, a key distinction for vectorial functions.
@@
Our work unifies these two lines of research by establishing the connection between symmetries and orbit enumeration for vectorial functions under EA-equivalence. We prove that almost all vectorial functions have trivial EA-stabilizers, extending the symmetry-breaking results of Clausen and Clote-Kranakis to the vectorial EA setting. As an immediate consequence via the orbit-stabilizer theorem, we recover the asymptotic orbit counts of Hou [5] and Lu et al. [6] with an elementary probabilistic proof, and reveal the underlying structural reason: the naive estimate (total number of functions divided by the group size) is asymptotically exact because stabilizers are trivial with high probability.
@@ 3.2 Probabilistic Analysis of EA-Stabilizers
This probability decays super-exponentially in $n$, demonstrating that vectorial Boolean functions with nontrivial EA-stabilizers are exponentially rare. In other words, cryptographic S-boxes generically possess no EA-symmetries: the overwhelming majority of functions in $\mathcal{F}$ have trivial stabilizer groups, and only an exponentially negligible fraction exhibit any symmetry under extended-affine transformations.
@@
Consequently, the contribution of functions with nontrivial EA-symmetries to the total count is exponentially negligible: their number is exponentially smaller than the full space. This observation is particularly relevant for structured searches in the design of cryptographic S-boxes.
@@
The upper bounds in Corollary 3.7 follow directly from the orbit-stabilizer theorem and hold for both EA- and CCZ-equivalence over arbitrary finite fields. However, using our result on trivial stabilizers (Theorem 3.4), we can show that the EA upper bound is asymptotically tight:
@@
Combining both contributions with the upper bound from Corollary 3.7 yields
@@ 4 Conclusion
    Classification: The number of EA-equivalence classes is asymptotically equal to the naive estimate (the total number of functions divided by the size of the EA-group) with vanishing relative error. This shows that the EA group action is asymptotically free: almost every function admits no nontrivial symmetries. For cryptographic applications, this means that the vast majority of vectorial functions lie in distinct EA-classes, and functions with nontrivial symmetries form an exponentially small subset.

### v2

# Almost All Vectorial Functions Have Trivial Extended-Affine Stabilizers

@@ Abstract
We prove that asymptotically almost all vectorial functions over finite fields have trivial extended-affine stabilizers. As a consequence, the number of EA-equivalence classes is asymptotically equal to the naive estimate, namely the total number of functions divided by the size of the EA-group, with vanishing relative error. Furthermore, we derive upper bounds on collision probabilities for both extended-affine and CCZ equivalences. For EA-equivalence, we leverage the trivial-stabilizer result to establish a matching lower bound, yielding a tight asymptotic formula that shows two independently sampled functions are EA-equivalent with super-exponentially small probability. The results validate random sampling strategies for cryptographic primitive design and show that functions with nontrivial EA-stabilizers form an exponentially rare subset.
Keywords: Boolean functions, extended-affine equivalence, EA-stabilizer, group action, orbit counting, asymptotic enumeration
@@ 1 Introduction
A substantial body of recent work has focused on the algorithmic and structural aspects of EA- and CCZ-equivalence. Canteaut and Perrin [3] showed that CCZ-equivalence can be decomposed into EA-equivalence and an operation called twisting, and derived upper bounds on the number of EA-equivalence classes within a single CCZ-equivalence class. Budaghyan, Calderini, and Villa [1] introduced sufficient conditions for determining when CCZ-equivalence is strictly more general than EA-equivalence combined with taking inverses of permutations. On the algorithmic side, Kaleyski [9] surveyed invariants for EA- and CCZ-equivalence of APN and AB functions and evaluated their effectiveness in distinguishing inequivalent instances. Canteaut, Couvreur, and Perrin [2] presented an efficient algorithm for recovering the EA-equivalence transformation between quadratic functions, along with a new invariant based on the ortho-derivative. These works address the problem of testing whether two given functions are EA-equivalent.
@@
Both questions are intimately connected to the structure of EA-stabilizer groups. The EA-stabilizer of a function $F$ is the set of all pairs of affine permutations $(A_{\rm in},A_{\rm out})$ satisfying $A_{\rm out}\circ F\circ A_{\rm in}=F$; it measures how much “self-equivalence” the function possesses. By the orbit-stabilizer theorem, the size of each equivalence class is inversely proportional to the stabilizer size, which directly impacts both the total number of classes and the collision probability. However, explicit computation of stabilizers is infeasible except for very small dimensions, as the function space grows doubly exponentially with the input dimension.
@@ 1.1 Our Contributions
Collectively, these results demonstrate that functions with nontrivial EA-stabilizers form an exponentially rare subset. This complements existing approaches focused on specific cryptographic properties and provides a theoretical foundation for understanding the prevalence of nontrivial stabilizers in the space of vectorial functions.
@@ 1.2 Related Work
Previous work on stabilizers of Boolean functions has primarily focused on simpler group actions. Clausen [6] proved that almost all Boolean functions (not vectorial) have trivial stabilizers under the action of $\mathrm{GL}(n,2)$. Clote and Kranakis [7] established similar results for invariance groups under variable permutations. However, these results apply only to Boolean functions $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ and do not address EA-equivalence, which involves affine transformations on both domain and range, a key distinction for vectorial functions.
@@
Our work unifies these two lines of research by establishing the connection between stabilizers and orbit enumeration for vectorial functions under EA-equivalence. We prove that almost all vectorial functions have trivial EA-stabilizers, extending the trivial-stabilizer results of Clausen and Clote-Kranakis to the vectorial EA setting. As an immediate consequence via the orbit-stabilizer theorem, we recover the asymptotic orbit counts of Hou [8] and Lu et al. [10] with an elementary probabilistic proof, and reveal the underlying structural reason: the naive estimate (total number of functions divided by the group size) is asymptotically exact because stabilizers are trivial with high probability.
@@ 3.2 Probabilistic Analysis of EA-Stabilizers
This probability decays super-exponentially in $n$, demonstrating that vectorial Boolean functions with nontrivial EA-stabilizers are exponentially rare. In other words, cryptographic S-boxes generically possess trivial EA-stabilizers: the overwhelming majority of functions in $\mathcal{F}$ have trivial stabilizer groups, and only an exponentially negligible fraction admit any nontrivial element in their EA-stabilizer.
Example 3.5 (Concrete bounds for cryptographic parameters).
We evaluate the bound from Theorem 3.4 for parameters commonly used in symmetric cryptography ($q=2$, $c=3/4$).
  1. 1.
    44-bit S-boxes ($n=m=4$): The group size satisfies $\log_{2}|\Gamma|=\log_{2}|\mathrm{AGL}(4,2)|^{2}=2(4+\log_{2}|\mathrm{GL}(4,2)|)=2(4+\log_{2}20160)<2\cdot 19=38$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 4\cdot 16=16$. Hence
    $\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{38-16}=2^{22},$
    which does not yield a nontrivial bound. This is expected: the function space $|\mathcal{F}|=2^{64}$ is not large enough relative to the group size for asymptotic freeness to manifest.
  2. 2.
    88-bit S-boxes ($n=m=8$): We have $\log_{2}|\Gamma|\leq 2(8^{2}+8)=144$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 8\cdot 256=512$. Hence
    $\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{144-512}=2^{-368}.$
    This shows that a uniformly random $8$-bit S-box has a nontrivial EA-stabilizer with probability at most $2^{-368}$, an astronomically small quantity. The collision probability for EA-equivalence is similarly bounded by $2^{-mq^{n}+O(n^{2}+m^{2})}=2^{-2048+O(128)}\approx 2^{-1900}$.
Thus, for dimensions $n\geq 8$ (the regime relevant to modern block ciphers such as AES), the trivial-stabilizer property holds with overwhelming probability.
@@
Consequently, the contribution of functions with nontrivial EA-stabilizers to the total count is exponentially negligible: their number is exponentially smaller than the full space. This observation is particularly relevant for structured searches in the design of cryptographic S-boxes.
@@
The upper bounds in Proposition 3.8 follow directly from the orbit-stabilizer theorem and hold for both EA- and CCZ-equivalence over arbitrary finite fields. However, using our result on trivial stabilizers (Theorem 3.4), we can show that the EA upper bound is asymptotically tight:
@@
Combining both contributions with the upper bound from Proposition 3.8 yields
@@ 4 Conclusion
    Classification: The number of EA-equivalence classes is asymptotically equal to the naive estimate (the total number of functions divided by the size of the EA-group) with vanishing relative error. This shows that the EA group action is asymptotically free: almost every function has a trivial EA-stabilizer. For cryptographic applications, this means that the vast majority of vectorial functions lie in distinct EA-classes, and functions with nontrivial EA-stabilizers form an exponentially small subset.